# Weak Proximal Newton Oracles for Composite Convex Optimization

### v1

@@ Abstract
Second-order methods are of great importance for composite convex optimization problems due to their local super-linear convergence rates (under appropriate assumptions). However, the presence of even a simple nonsmooth function in the model most often renders the subproblems in proximal Newton methods computationally-difficult to solve in high-dimension. We introduce a novel approach based on a weak proximal Newton oracle (WPNO), which only requires to solve such subproblems to accuracy that is comparable to that of the optimal solution of the global problem, while maintaining local super-linear convergence under standard assumptions. We show that when the optimal solution of the global problem admits some sparse structure, a WPNO could be implemented very efficiently using specialized first-order methods.
@@ 1 Introduction
[…]
@@
$\forall{\mathbf{x}}\in\textrm{dom}(\mathcal{R}):\quad\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|^{2}\leq\frac{2}{\alpha}\left({F({\mathbf{x}})-F^{*}}\right).$
In this paper we will be interested in efficient second-order methods for Problem (1) with local super-linear convergence rates (under either Assumption 1 or 2), which could highly beneficial, at least at a certain proximity of the optimal solution, when Problem (1) is ill-conditioned for first-order methods in the sense that the first-order condition number $\beta/\alpha$ is very large, where $\beta$ denotes the Lipschitz continuity parameter of $\nabla{}f$.
The literature on second-order methods for convex optimization is of course huge, and we do not presume to thoroughly survey it here. Instead we refer the interested reader to [14, 21, 20, 7, 16, 22, 5] and references therein.
Standard Newton methods, when applied to Problem (1) often require on each iteration $t$ to solve a subproblem of the form:
@@
[…]
@@
Problem (2) is most often difficult to solve even when $\mathcal{R}$ is quite simple. For instance, already when $f$ is non-linear and $\mathcal{R}$ is the indicator function for an $\ell_{1}$ ball in $\mathbb{R}^{n}$, Problem (2) does not admit a closed-form solution and requires the use of iterative optimization methods, e.g., first-order methods, to solve it to sufficient accuracy, within each iteration of the Newton method.
@@
Unfortunately, solving Problem (2) using standard first-order methods can often defeat the purpose of using a second-order method. For instance, even an optimal first-order method with an accelerated linear convergence-rate ,e.g., [18, 3, 17], will require in worst-case $\sqrt{\beta/\alpha}$ number of gradient computations of $\Psi_{t}(\cdot)$ and prox computations (e.g., Euclidean projections) w.r.t. $\mathcal{R}$. Computing the gradient of $\Psi_{t}$ requires in general dense matrix-vector product computations (multiplication with Hessian matrix), and in case $\mathcal{R}$ is for instance the indicator function for a nuclear norm ball of matrices or a non-trivial polytope, computing the prox operator may also be computationally difficult. Thus, such methods will introduce computationally-intensive operations into the Newton method whose number scales with the first-order condition number $\beta/\alpha$, which may ultimately limit the benefit of using second-order information.
@@
In case the bottleneck in solving (2) via proximal gradient methods is the complex structure of $\mathcal{R}$, some works such as [6, 15] have considered using the Frank-Wolfe (aka conditional gradient) method, which only requires a linear minimization oracle w.r.t. $\textrm{dom}(\mathcal{R})$. However, the slow convergence rate of the Frank-Wolfe method, which is typically $1/t$, may again result in an overall prohibitive runtime, in particular when the overall desired accuracy is medium-high. In case $\mathcal{R}$ is the indicator function for a polytope, some variants of Frank-Wolfe enjoy a linear convergence rate, however this comes, in worst-case, with a multiplicative factor that scales at least linearly with both the first-order condition number of $f$ and the ambient dimension of the problem [6]. We note that while some Frank-Wolfe variants for polytopes can replace the explicit dependence on the ambient dimension with the dimension of the optimal face (containing the optimal solution), which can indeed be low-dimensional in certain settings [11, 1], even if ${\mathbf{x}}^{*}$ indeed lies on a low-dimensional face, there is in general no guarantee that the solutions to the subproblems (2) will also lie on a low-dimensional face.
We shall also mention the classical work [4] that considered the case in which $\mathcal{R}$ is the indicator function for the non-negative orthant in $\mathbb{R}^{n}$, and considered a method that is based on a considerably simplified version of subproblem (2), namely replacing the computation of the projection onto the non-negative orthant w.r.t. the norm induced by the Hessian $\nabla^{2}f({\mathbf{x}}_{t})$, with the standard Euclidean norm. [4] established a local super-linear convergence for this simplified method, however it holds only under an additional strong assumption known as strict complementarity which we do not make here (see Proposition 4 in [4]).
@@
In this work we develop a new principled approach towards making proximal Newton methods more efficient to implement while maintaining local super-linear convergence rates. Our approach will be mostly of interest in case the optimal solution ${\mathbf{x}}^{*}$ admits a certain sparse structure. Indeed in many cases, for example when $\mathcal{R}$ is the indicator function for an $\ell_{1}$ ball or matrix nuclear norm ball, or even a convex and compact polytope, we may expect that ${\mathbf{x}}^{*}$ is indeed sparse, may it be entry-wise sparsity for the $\ell_{1}$ ball case, low rank for the matrix nuclear norm ball case, or that it lies on a low-dimensional face in the polytope case.
@@
Our approach is based on two main observations. First, we observe that in order to obtain local super-linear convergence (in function value and under either Assumption 1 or Assumption 2), subproblem (2) (or a regularized variant of it) need not be solved to optimality (or even with very good approximation), but only to value that is comparable with that of ${\mathbf{x}}^{*}$ — the optimal solution to our original Problem (1). We refer to these as weak proximal Newton oracles (Definition 1 in the sequel). Second, we show that specialized first-order methods, when applied to subproblems such as (2), but require to solve it only to the level of error achieved by ${\mathbf{x}}^{*}$, can leverage the sparse structure of ${\mathbf{x}}^{*}$ to solve these with better complexity than standard first-order methods. This is despite the fact that the optimal solutions to subproblem (2) need not inherit the sparse structure of ${\mathbf{x}}^{*}$ in general. The specialized first-order methods considered include those based on first-order weak proximal oracles, recently introduced in [9, 10], and decomposition-invariant conditional gradient methods [11, 1].
@@
The rest of this paper is structured as follows: In section 2 we define our main object of interest — weak proximal Newton oracles (WPNO), and establish their local super-linear convergence. In Sections 3 and 4 we discuss efficient implementations based on first-order weak proximal oracles, and conditional gradient methods, respectively. Finally, in Section 5 we bring some preliminary empirical evidence to demonstrate the plausibility of our approach.
@@ 2 Weak Proximal Newton Oracles
Let us comment why we consider both regularized and unregularized WPNOs. As we shall see in Theorem 1, the use of the regularized WRPNO will lead to a preferable super-linear convergence rate, however it will mostly be interesting when assuming Assumption 2 holds. This is because under Assumption 2, the function $\phi_{{\mathbf{x}},\eta}^{r}$ admits a bounded Lipchitz parameter for the gradient of its smooth part over $\textrm{dom}(\mathcal{R})$ (since $\textrm{dom}(\mathcal{R})$ itself is bounded), which is compatible with the use of first-order methods to implement the WRPNO. When only Assumption 1 holds, we can no longer assume the smooth part of $\phi_{{\mathbf{x}},\eta}^{r}$ has bounded Lipchitz parameter for its gradient (due to the cubic term), and instead we rely on its unregularized version $\phi_{{\mathbf{x}},\eta}$, which is in particular quadratic, but which will lead to a slightly inferior (in terms of constants) super-linear convergence rate.
@@
Lemm 1 leads to the following inequality which will be of use.
@@
The following lemma bounds the distance of the output of a WPNO (Definition 2) from the optimal solution ${\mathbf{x}}^{*}$.
@@
Proof.
Define the function:
We have that,
=$\underset{(b)}{=}\phi_{{\mathbf{x}},\eta}({\mathbf{v}})-\phi_{{\mathbf{x}},\eta}({\mathbf{x}}^{*})-\frac{\eta}{2}\|{{\mathbf{v}}-{\mathbf{x}}}\|_{{\mathbf{x}}}^{2}+\frac{\eta}{2}\|{{\mathbf{x}}^{*}-{\mathbf{x}}}\|_{{\mathbf{x}}}^{2}$
$\underset{(c)}{\leq}\eta^{2}C_{\mathcal{A}}\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|^{3}$
Note that,
Plugging into the previous inequality, we have that
Combining Eq. (7) and Eq. (2) we obtain,
@@ 2.1 Local super-linear convergence
$\forall t\geq 1:\qquad{\mathbf{x}}_{t+1}\leftarrow\left\{\begin{array}[]{ll}\mathcal{A}({\mathbf{x}}_{t+1},1)&\mbox{if $F(\mathcal{A}({\mathbf{x}}_{t+1},1))<F({\mathbf{x}}_{t})$};\\ {\mathbf{x}}_{t}&else.\end{array}\right.$ (12)
@@
In both cases of the theorem, if $C_{\mathcal{A}}=O(\beta_{2})$, then the local super-linear convergence rate is of the form $F({\mathbf{x}}_{t+1})-F^{*}=O\left({\left({F({\mathbf{x}}_{t})-F^{*}}\right)^{3/2}}\right)$. In particular, the rate in (13) is the same, up to a universal constant, as the rate obtained in [20] (Theorem 5) for the exact cubic-regularized Newton method in the unconstrained case (i.e., $\mathcal{R}\equiv 0$).
@@ 3 Implementing WPNOs via First-Order Weak Proximal Oracles
In this section we discuss the efficient implementation of WPNOs (Definition 1) using specialized first-order methods based on first-order weak proximal oracles [9, 10]. In particular, we shall derive novel convergence results for these first-order methods that allow to leverage sparse structure of ${\mathbf{x}}^{*}$, even when applied to optimization problems whose optimal solutions do not necessarily admit such sparse structure.
@@
In Section 3.1 we discuss in detail cases of interest in which a first-order weak proximal oracle could be implemented very efficiently be leveraging sparse structures in ${\mathbf{x}}^{*}$.
@@
If Assumption 1 holds and $\phi_{t}=\phi_{{\mathbf{x}}_{t},1}$, then the sequence $({\mathbf{y}}_{i})_{i\geq 1}$ produced by Algorithm 1 with a fixed step-size $\lambda_{i}=\alpha/\tilde{\beta}$ ($\alpha$ as in Assumption 1) satisfies for all $i\geq 1$:
@@
If Assumption 2 holds and $\phi_{t}=\phi^{r}_{{\mathbf{x}}_{t},1}$, then using a fixed step-size $\lambda_{i}=\alpha/\tilde{\beta}$ ( $\alpha$ as in Assumption 2) satisfies that for all $i\geq 1$:
@@
Thus, we can conclude that in both scenarios (Assumption 1 or Assumption 2), by running Algorithm 1 for sufficiently-many iterations ($\tilde{O}(\tilde{\beta}/\alpha)$, where $\tilde{O}(\cdot)$ suppresses logarithmic factors), we can implement a W(R)PNO with parameter $C_{\mathcal{A}}=O(\beta_{2})$.

### v2

@@ Abstract
Second-order methods are of great importance for composite convex optimization problems due to their local super-linear convergence rates (under appropriate assumptions). However, the presence of even a simple nonsmooth function in the model most often renders the subproblems in proximal Newton methods computationally difficult to solve in high dimensions. We introduce a novel approach based on a weak proximal Newton oracle (WPNO), which only requires solving such subproblems to accuracy that is comparable to that of the optimal solution of the global problem, while maintaining local super-linear convergence under standard assumptions. Mainly, unlike classical inexact proximal Newton schemes, the complexity of our WPNO is not tied to (approximately) minimizing each subproblem; instead, we establish that when the optimal solution of the global problem admits a sparse structure, the inner subproblem can be solved by specialized first-order methods whose cost scales directly with the sparsity of this solution rather than with the ambient dimension.
@@ 1 Introduction
where $\mathbb{E}$ is a Euclidean vector space of dimension $n$, $\mathcal{R}:\mathbb{E}\rightarrow(-\infty,\infty]$ is a proper closed convex function, and $f(\cdot)$ is convex and twice continuously differentiable over $\textrm{dom}(\mathcal{R})$. We let $\beta_{2}$ denote the Lipschitz continuity parameter of the Hessian of $f$, i.e., $\|{\nabla^{2}f({\mathbf{x}})-\nabla^{2}f({\mathbf{y}})}\|\leq\beta_{2}\|{{\mathbf{x}}-{\mathbf{y}}}\|$ for all ${\mathbf{x}},{\mathbf{y}}\in\textrm{dom}(\mathcal{R})$, where $\|{\cdot}\|$ denotes the Euclidean norm. Throughout we shall assume that Problem (1) is bounded from below and we shall denote the optimal value by $F^{*}$. Throughout we shall also assume either one of the following two standard assumptions holds true.
@@
Note Assumption 1 implies that there is a unique minimizer $\mathbf{x}$$^{*}$ to Problem (1).
@@
$\forall{\mathbf{x}}\in\textrm{dom}(\mathcal{R}):\quad\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|^{2}\leq\frac{2}{\alpha}\left({F({\mathbf{x}})-F^{*}}\right),$
where $\mathbf{x}$$^{*}$ is the unique minimizer of Problem (1)111since our derivations will be often interesting when $\mathcal{R}$ is a sparsity-promoting function, and when the optimal solutions are indeed sparse in the appropriate sense, the assumption here that $\mathbf{x}$$^{*}$ is unique is quite plausible. Nevertheless, this assumption is virtually without loss of generality, as the extension of the analyzes to handle multiple solutions is quite straightforward.
In this paper we will be interested in efficient second-order methods for Problem (1) with local super-linear convergence rates (under either Assumption 1 or 2), which could be highly beneficial, at least at a certain proximity of the optimal solution, when Problem (1) is ill-conditioned for first-order methods in the sense that the first-order condition number $\beta/\alpha$ is very large, where $\beta$ denotes the Lipschitz continuity parameter of $\nabla{}f$.
The literature on second-order methods for convex optimization is of course huge, and we do not presume to thoroughly survey it here. Instead we refer the interested reader to [16, 24, 23, 9, 19, 6, 17, 25, 5] and references therein.
Standard Newton methods, when applied to Problem (1), often require on each iteration $t$ to solve a subproblem of the form:
@@
where ${\mathbf{x}}_{t}$ is the current iterate, $\eta>0$ is a step-size, and $\langle{\cdot,\cdot}\rangle$ is the standard inner-product, see for instance [16, 24, 9, 25]. Some methods consider a cubic-regularized variant of (2), see for instance [23, 22].
@@
Problem (2) is most often difficult to solve even when $\mathcal{R}$ is quite simple. For instance, already when $f$ is non-linear and $\mathcal{R}$ is the indicator function for an $\ell_{1}$ ball in $\mathbb{R}^{n}$, Problem (2) does not admit a closed-form solution and requires the use of iterative optimization methods, e.g., first-order methods, to solve it to sufficient accuracy, within each iteration of the Newton method. Indeed there is a significant body of work on inexact proximal Newton methods that maintain a local super-linear convergence rate, provided that Problem (2) is solved to error that is driven sufficiently fast (with the Newton iterations counter $t$) to zero, see for instance [16, 6, 17] and references therein.
@@
Unfortunately, solving Problem (2) within inexact Newton methods using standard first-order methods can often defeat the purpose of using a second-order method. For instance, even an optimal first-order method with an accelerated linear convergence rate ,e.g., [21, 3, 20], will require in worst-case number of gradient computations of (the smooth part of) $\Psi_{t}(\cdot)$ and prox computations (e.g., Euclidean projections) w.r.t. $\mathcal{R}$ that scale with $\sqrt{\beta/\alpha}$. Computing the gradient of $\Psi_{t}$ requires in general a dense matrix-vector product computation (multiplication with Hessian matrix), and in case $\mathcal{R}$ is for instance the indicator function for a nuclear norm ball of matrices or a non-trivial polytope, computing the prox operator may also be computationally difficult. Thus, the inner complexity bounds of such inexact Newton methods are typically expressed in terms of the condition number of $\Psi_{t}$ and the ambient dimension, without taking advantage of any sparsity, low-rankness, or low-face structure of the global optimum $\mathbf{x}$$^{*}$.
@@
In case the bottleneck in solving (2) via proximal gradient methods is the complex structure of $\mathcal{R}$, some works such as [7, 18] have considered using the Frank-Wolfe (aka conditional gradient) method, which only requires a linear minimization oracle w.r.t. $\textrm{dom}(\mathcal{R})$. However, the slow convergence rate of the Frank-Wolfe method, which is typically $1/k$, may again result in an overall prohibitive runtime, in particular when the overall desired accuracy is medium-high. In case $\mathcal{R}$ is the indicator function for a polytope, some variants of Frank-Wolfe enjoy a linear convergence rate, however this comes, in worst-case, with a multiplicative factor that scales at least linearly with the product of the first-order condition number of $\Psi_{t}$ and the ambient dimension [7]. We note that while some Frank-Wolfe variants for polytopes can replace the explicit dependence on the ambient dimension with the dimension of the optimal face (containing the optimal solution), which can indeed be low-dimensional in certain settings [13, 1], even if $\mathbf{x}$$^{*}$ indeed lies on a low-dimensional face, there is in general no guarantee that the solutions to the subproblems (2) will also lie on a low-dimensional face.
The above discussions lead us to the following conceptual question:
In case the optimal solution to Problem (1) admits some known sparse structure, can it be leveraged towards provable more efficient implementations of proximal Newton methods, without sacrificing the local super-linear convergence?
@@
In this work we develop a new principled approach towards making proximal Newton methods more efficient to implement while maintaining local super-linear convergence rates. Indeed our approach will be mostly of interest in case the optimal solution $\mathbf{x}$$^{*}$ admits a certain sparse structure. For example, when $\mathcal{R}$ is the indicator function for an $\ell_{1}$ ball or a matrix nuclear norm ball, or even a polytope, we may often expect that $\mathbf{x}$$^{*}$ is indeed sparse, may it be entry-wise sparsity for the $\ell_{1}$ ball case, low rank for the matrix nuclear norm ball case, or that it lies on a low-dimensional face in the polytope case.
@@
Our approach is based on two key observations. First, we observe that in order to obtain local super-linear convergence (in function value and under either Assumption 1 or Assumption 2), subproblem (2) (or a regularized variant of it) need not be solved to optimality (or even with very good accuracy as in classical inexact proximal Newton analyzes [16, 6, 17]), but only to value that is comparable with that of $\mathbf{x}$$^{*}$ — the optimal solution to our original Problem (1). We refer to these as weak proximal Newton oracles (Definition 1 in the sequel). Second, and perhaps surprisingly, we are able to show that specialized first-order methods, when applied to subproblems such as (2), but require to solve it only to the level of error achieved by $\mathbf{x}$$^{*}$, can solve these with complexity that scales directly with the sparsity of $\mathbf{x}$$^{*}$ and not with the ambient dimension as in standard first-order methods, despite the fact that in general the optimal solutions to subproblem (2) need not inherit the sparse structure of $\mathbf{x}$$^{*}$. For instance, if $\mathcal{R}$ is the indicator for an $\ell_{1}$-ball and $\mathbf{x}$$^{*}$ is (as often expected in this case) entry-wise sparse, then our solvers for subproblem (2) will require to compute the product of the Hessian with only a sparse vector (with same level of sparsity as assumed for $\mathbf{x}$$^{*}$), instead of a dense vector as in standard gradient methods. In case $\mathcal{R}$ is the indicator for a matrix nuclear norm ball and $\mathbf{x}$$^{*}$ is low-rank (again, as often expected when using such $\mathcal{R}$), our solvers will require to compute only a low-rank singular value decomposition (SVD) (with same rank as the one assumed for $\mathbf{x}$$^{*}$), instead of a full-rank SVD required for the projection step onto the nuclear norm ball in standard gradient methods, which is computationally prohibitive in high dimensions. In case $\mathcal{R}$ is a polytope suitable for Frank-Wolfe/conditional gradient methods (i.e., implementing the linear minimization oracle is efficient), our conditional gradient-based solver will converge linearly and with number of iterations that scales with the dimension of the optimal face containing $\mathbf{x}$$^{*}$ and not with the ambient dimension as in standard linearly-converging conditional gradient methods. The specialized first-order methods which we develop are based on first-order weak proximal oracles, recently introduced in [11, 12], and decomposition-invariant conditional gradient methods [13, 1].
@@
1.1 Organization of this paper
In section 2 we define our main object of interest — weak proximal Newton oracles (WPNO), and establish their local super-linear convergence. In Section 3 we derive efficient implementations of our WPNOs based on first-order weak proximal oracles. In this section we also discuss concrete examples where our approach may significantly improve over standard methods. In section 4 we discuss efficient implementations of our WPNOs based on a state-of-the-art conditional gradient method. Finally, in Section 5 we bring empirical evidence to demonstrate the potential practical benefits of our results.
@@ 2 Weak Proximal Newton Oracles
In other words and less formally: A WPNO solves the same type of subproblem as in the proximal Newton method (e.g., [16]), but does so only to an approximation error competitive with the global optimal solution $\mathbf{x}$$^{*}$, i.e., up to an additive error scaling with $\|{{\mathbf{x}}-{\mathbf{x}}^{*}}\|^{3}$ (where $\mathbf{x}$} is the current point) from $\mathbf{x}$$^{*}$. A WRPNO does the same but w.r.t. the type of subproblem appearing in the cubic-regularized Newton method (e.g. [23]).
Let us comment why we consider both regularized and unregularized WPNOs. As we shall see in Theorem 1, the use of the regularized WRPNO will lead to a preferable super-linear convergence rate, however it will mostly be interesting when Assumption 2 holds. This is because under Assumption 2, the function $\phi_{{\mathbf{x}},\eta}^{r}$ admits a bounded Lipschitz parameter for the gradient of its smooth part over $\textrm{dom}(\mathcal{R})$ (since $\textrm{dom}(\mathcal{R})$ itself is bounded), which is compatible with the use of first-order methods to implement the WRPNO. When only Assumption 1 holds, we can no longer assume the smooth part of $\phi_{{\mathbf{x}},\eta}^{r}$ has bounded Lipschitz parameter for its gradient (due to the cubic term), and instead we rely on its unregularized version $\phi_{{\mathbf{x}},\eta}$, which is in particular quadratic, but which will lead to a slightly inferior (in terms of constants) super-linear convergence rate.
@@
We now present several lemmas that will be important for our derivations.
@@
Lemma 1 leads to the following inequality which will be of use.
@@
The following lemma bounds the distance of the output of a WPNO (Definition 2) from the optimal solution $\mathbf{x}$$^{*}$. For clarity of presentation, we defer the proof to the appendix.
@@ 2.1 Local super-linear convergence
$\forall t\geq 1:\qquad{\mathbf{x}}_{t+1}\leftarrow\left\{\begin{array}[]{ll}\mathcal{A}({\mathbf{x}}_{t},1)&\mbox{if $F(\mathcal{A}({\mathbf{x}}_{t},1))<F({\mathbf{x}}_{t})$};\\ {\mathbf{x}}_{t}&else.\end{array}\right$. (9)
@@
In both cases of the theorem, if $C_{\mathcal{A}}=O(\beta_{2})$, then the local super-linear convergence rate is of the form $F({\mathbf{x}}_{t+1})-F^{*}=O\left({\left({F({\mathbf{x}}_{t})-F^{*}}\right)^{3/2}}\right)$, provided that the intialization point ${\mathbf{x}}_{1}$ is such that the initial error $F({\mathbf{x}}_{1})-F^{*}$ is already sufficiently small. In particular, the rate in (10) is the same, up to a universal constant, as the rate obtained in [23] (Theorem 5) for the exact cubic-regularized Newton method in the unconstrained case (i.e., $\mathcal{R}\equiv 0$).
@@ 3 Implementing WPNOs via First-Order Weak Proximal Oracles
In this section we discuss the efficient implementation of WPNOs (Definition 1) using specialized first-order methods based on first-order weak proximal oracles [11, 12]. In particular, we shall derive novel convergence results for these first-order methods that allow to leverage the sparse structure of $\mathbf{x}$$^{*}$, even when applied to optimization problems whose optimal solutions do not necessarily admit such sparse structure.
@@
In Section 3.1 we discuss in detail cases of interest in which a first-order weak proximal oracle could be implemented very efficiently by leveraging sparse structures in $\mathbf{x}$$^{*}$.
@@
If Assumption 1 holds and $\phi_{t}=\phi_{{\mathbf{x}}_{t},1}$, then the sequence $({\mathbf{y}}_{i})_{i\geq 1}$ produced by Algorithm 1 with a fixed step-size $\lambda_{i}=\frac{\alpha}{\tilde{\beta}}$ ($\alpha$ as in Assumption 1) satisfies for all $i\geq 1$:
@@
If Assumption 2 holds and $\phi_{t}=\phi^{r}_{{\mathbf{x}}_{t},1}$, then using a fixed step-size $\lambda_{i}=\frac{\alpha}{2\tilde{\beta}}$ ( $\alpha$ as in Assumption 2) satisfies that for all $i\geq 1$:
@@
[…]